OpenStudy (anonymous):

Find the sums and show work 1+2+3+ ... +498+499+500

OpenStudy (anonymous):

n(n+1)/2 i think with n being the number of numbers

jhonyy9 (jhonyy9):

write those all hence in this way 1+499=500 2+498=500 3+497=500

jhonyy9 (jhonyy9):

at finaly you need numerating how many from 500 there are ?

jhonyy9 (jhonyy9):

multiply 500 with this get number and will get the result

OpenStudy (anonymous):

What he said. You have 250 sets of 500 by making 500 though 499 =1, 498 +2 etc etc but you also have 500 + 0 so you have 251 sets of 500. 251*500= 125500.